# Confabulation dynamics in a reservoir computer: Filling in the gaps with untrained attractors

### v2

@@ Appendix C Additional reconstruction routes for $\pazocal{L}$
In this Appendix we illustrate some less common routes which we find the RC takes to reconstruct $\pazocal{L}$ in task (i). The routes of reconstruction we illustrate here correspond to the classification of the RC’s output at different $\rho$ values for $\textbf{M}_{19}$ and $\textbf{M}_{27}$ as shown in Fig. 7. We use the same method that was used in Fig. 9 to track the changes in the dynamics of the reconstructed attractor. For both $\textbf{M}_{19}$ and $\textbf{M}_{27}$ we encounter no UAs for the values of $\rho$ where the tracking is performed. More specifically, we plot $x_{3}^{max}$ (local maxima of the $x_{3}$ variable) and the algorithm’s classification at specified values of $\rho$ in Figs. 25 and 27. In Figs. 26 and 28 we illustrate the RC’s reconstruction of $\pazocal{L}$ at key points indicated by the inward ticks in Figs. 25 and 27.
C.0.1 Reconstruction routes found from $\textbf{M}_{19}$
The coloured bars and bifurcation diagram in Fig. 25 show that while the RC’s reconstruction of $\pazocal{L}$ is reasonably good for $\rho=0.6$, the reconstruction worsens before it improves for increasing $\rho$ from $0.6$ to $0.87$. Within this range of $\rho$ values the poor reconstructions of $\pazocal{L}$, $\tilde{\pazocal{L}}$, undergo several changes between chaotic and periodic dynamics. Fig. 28 highlights some of these changes.
The first major change in the RC’s reconstruction of $\pazocal{L}$ occurs nearby $\rho=0.625$, the transition from $\hat{\pazocal{L}}$ to $\tilde{\pazocal{L}}$ as shown in Figs. 28 (a) and (b) involves a break in the symmetry of the reconstructed attractor. However what is more interesting about the dynamics of $\tilde{\pazocal{L}}$ for this M, and the changes it undergoes for increasing $\rho$ as shown in panels (b) to (i), is that it provides a much clearer indication of the mechanisms that give rise to this interesting break in the symmetry of the RC’s dynamics. From looking at these changes in the dynamics of $\tilde{\pazocal{L}}$, the prolonged winding motion nearby the point $(x_{1},x_{3})=(8,20)$ followed by the much larger orbit strongly resembles a saddle-focus homoclinic orbit, indicating that a Shilnikov bifurcation occurs. Figs. 25 and 28 show that this attractor exhibits periodic dynamics for a wide range of $\rho$ values and also undergoes a period-doubling route to chaos nearby $\rho=0.675$. While this attractor alternates between chaotic and periodic dynamics, what is most interesting here is that the chaotic dynamics are largely confined to the portion of the trajectory that revolves around the saddle-focus and while that the homoclinic orbit portion of the trajectory can also be chaotic (like in (b)) it is mainly periodic (like in (c)–(h)) depending on the value of $\rho$. In other words, in (b) there are multiple routes in and out of the saddle-focus portion of the trajectory on the attractor but in (c)–(h) there is typically only one or two routes in and out. For a small range of $\rho$ values nearby $0.785$ we see in Fig. 25 (i) that there is no homoclinic orbit component to $\tilde{\pazocal{L}}$. The growth of the saddle-focus portion of the attractor from (f)–(h) indicates that the transition from (h) to (i) may be due to a collision with a saddle. Increasing $\rho$ further we find that the reconstruction alternates between a more Lorenz-like chaotic attractor and the saddle-focus homoclinic orbit which also alternates between chaotic and periodic dynamics as shown in Figs. 25 (j)–(o). For $\rho>0.825$ we see that reasonably good reconstructions of $\pazocal{L}$, like in Fig. 25 (o), become more and more prevalent.
@@ C.0.1 Reconstruction routes found from $\textbf{M}_{19}$
C.0.2 Reconstruction routes found from $\textbf{M}_{27}$
The coloured bars and bifurcation diagram in Fig. 25 show that the RC’s reconstruction of $\pazocal{L}$ improves by increasing $\rho$, however there is by no means a gradual improvement from the poor reconstruction, $\tilde{\pazocal{L}}$, to the good reconstruction, $\hat{\pazocal{L}}$, as $\tilde{\pazocal{L}}$ undergoes a wide-range of changes from $\rho=0.6$ to $1.25$.
The first of the major changes in the dynamics of $\tilde{\pazocal{L}}$ occurs around $\rho=0.66$ where the Lorenz-like chaotic attractor in Fig. 26 (a) becomes a limit cycle in (b). What is interesting about this limit cycle is that there is a clear break in the symmetry of the RC’s dynamics, two round trips of the right-hand wing occur before a single trip around the left-hand wing. These periodic dynamics do not persist for many $\rho$ values as $\tilde{\pazocal{L}}$ becomes chaotic again near $\rho=0.67$. However, it is after the return to chaos that we see break in the symmetry of the RC’s dynamics becomes much more pronounced, the range of values taken by $x_{3}^{max}$ for the right-hand wing of $\tilde{\pazocal{L}}$ increases as $\rho$ is increased to $0.8633$. These changes in $\tilde{\pazocal{L}}$ are further highlighted by Figs. 26 (c)–(e), it is also worth noting in (e) that we indirectly observe the presence of a saddle point given the sharp bend in the trajectory nearby $(x_{1},x_{3})=(0,0)$ thereby indicating that the attractor is increasing in volume and approaching a saddle point. This may be responsible for the return to periodic dynamics in (f) and (g) that we see as $\rho$ is increased further. In contrast to the limit cycle shown previously in (b), we see in (f) and (g) there are more round trips on the left-hand wing. $\tilde{\pazocal{L}}$ alternates between chaotic and periodic dynamics for increasing $\rho$ and the range of values taken by $x_{3}^{max}$ for the right-hand wing of $\tilde{\pazocal{L}}$ also starts to decrease. The first instances of good reconstructions of $\pazocal{L}$, namely $\hat{\pazocal{L}}$, emerge near $\rho=1.215$ and the RC maintains a reasonably good reconstruction of $\pazocal{L}$ for $\rho>1.24$.

### v3

@@ Appendix D Additional reconstruction routes for $\pazocal{L}$
In this Appendix we illustrate some less common routes which we find the RC takes to reconstruct $\pazocal{L}$ in task (i). The routes of reconstruction we illustrate here correspond to the classification of the RC’s output at different $\rho$ values for $\textbf{M}_{19}$ and $\textbf{M}_{27}$ as shown in Fig. 8. We use the same method that was used in Fig. 9 to track the changes in the dynamics of the reconstructed attractor. For both $\textbf{M}_{19}$ and $\textbf{M}_{27}$ we encounter no UAs for the values of $\rho$ where the tracking is performed. More specifically, we plot $x_{3}^{max}$ (local maxima of the $x_{3}$ variable) and the algorithm’s classification at specified values of $\rho$ in Figs. 22 and 24. In Figs. 23 and 25 we illustrate the RC’s reconstruction of $\pazocal{L}$ at key points indicated by the inward ticks in Figs. 22 and 24.
D.0.1 Reconstruction routes found from $\textbf{M}_{19}$
@@
The coloured bars and bifurcation diagram in Fig. 22 show that while the RC’s reconstruction of $\pazocal{L}$ is reasonably good for $\rho=0.6$, the reconstruction worsens before it improves for increasing $\rho$ from 0.60.60.6 to 0.870.870.87. Within this range of $\rho$ values the poor reconstructions of $\pazocal{L}$, $\tilde{\pazocal{L}}$, undergo several changes between chaotic and periodic dynamics. Fig. 25 highlights some of these changes.
The first major change in the RC’s reconstruction of $\pazocal{L}$ occurs nearby $\rho=0.625$, the transition from $\hat{\pazocal{L}}$ to $\tilde{\pazocal{L}}$ as shown in Figs. 25 (a) and (b) involves a break in the symmetry of the reconstructed attractor. However what is more interesting about the dynamics of $\tilde{\pazocal{L}}$ for this M, and the changes it undergoes for increasing $\rho$ as shown in panels (b) to (i), is that it provides a much clearer indication of the mechanisms that give rise to this interesting break in the symmetry of the RC’s dynamics. From looking at these changes in the dynamics of $\tilde{\pazocal{L}}$, the prolonged winding motion nearby the point $(x_{1},x_{3})=(8,20)$ followed by the much larger orbit strongly resembles a saddle-focus homoclinic orbit, indicating that a Shilnikov bifurcation occurs. Figs. 22 and 25 show that this attractor exhibits periodic dynamics for a wide range of $\rho$ values and also undergoes a period-doubling route to chaos nearby $\rho=0.675$. While this attractor alternates between chaotic and periodic dynamics, what is most interesting here is that the chaotic dynamics are largely confined to the portion of the trajectory that revolves around the saddle-focus and while that the homoclinic orbit portion of the trajectory can also be chaotic (like in (b)) it is mainly periodic (like in (c)–(h)) depending on the value of $\rho$. In other words, in (b) there are multiple routes in and out of the saddle-focus portion of the trajectory on the attractor but in (c)–(h) there is typically only one or two routes in and out. For a small range of $\rho$ values nearby 0.7850.7850.785 we see in Fig. 22 (i) that there is no homoclinic orbit component to $\tilde{\pazocal{L}}$. The growth of the saddle-focus portion of the attractor from (f)–(h) indicates that the transition from (h) to (i) may be due to a collision with a saddle. Increasing $\rho$ further we find that the reconstruction alternates between a more Lorenz-like chaotic attractor and the saddle-focus homoclinic orbit which also alternates between chaotic and periodic dynamics as shown in Figs. 22 (j)–(o). For $\rho>0.825$ we see that reasonably good reconstructions of $\pazocal{L}$, like in Fig. 22 (o), become more and more prevalent.
D.0.2 Reconstruction routes found from $\textbf{M}_{27}$
@@
The coloured bars and bifurcation diagram in Fig. 22 show that the RC’s reconstruction of $\pazocal{L}$ improves by increasing $\rho$, however there is by no means a gradual improvement from the poor reconstruction, $\tilde{\pazocal{L}}$, to the good reconstruction, $\hat{\pazocal{L}}$, as $\tilde{\pazocal{L}}$ undergoes a wide-range of changes from $\rho=0.6$ to 1.251.251.25.
The first of the major changes in the dynamics of $\tilde{\pazocal{L}}$ occurs around $\rho=0.66$ where the Lorenz-like chaotic attractor in Fig. 23 (a) becomes a limit cycle in (b). What is interesting about this limit cycle is that there is a clear break in the symmetry of the RC’s dynamics, two round trips of the right-hand wing occur before a single trip around the left-hand wing. These periodic dynamics do not persist for many $\rho$ values as $\tilde{\pazocal{L}}$ becomes chaotic again near $\rho=0.67$. However, it is after the return to chaos that we see break in the symmetry of the RC’s dynamics becomes much more pronounced, the range of values taken by $x_{3}^{max}$ for the right-hand wing of $\tilde{\pazocal{L}}$ increases as $\rho$ is increased to 0.86330.86330.8633. These changes in $\tilde{\pazocal{L}}$ are further highlighted by Figs. 23 (c)–(e), it is also worth noting in (e) that we indirectly observe the presence of a saddle point given the sharp bend in the trajectory nearby $(x_{1},x_{3})=(0,0)$ thereby indicating that the attractor is increasing in volume and approaching a saddle point. This may be responsible for the return to periodic dynamics in (f) and (g) that we see as $\rho$ is increased further. In contrast to the limit cycle shown previously in (b), we see in (f) and (g) there are more round trips on the left-hand wing. $\tilde{\pazocal{L}}$ alternates between chaotic and periodic dynamics for increasing $\rho$ and the range of values taken by $x_{3}^{max}$ for the right-hand wing of $\tilde{\pazocal{L}}$ also starts to decrease. The first instances of good reconstructions of $\pazocal{L}$, namely $\hat{\pazocal{L}}$, emerge near $\rho=1.215$ and the RC maintains a reasonably good reconstruction of $\pazocal{L}$ for $\rho>1.24$.
Appendix E Examining the continuous transition between $\hat{\pazocal{L}}$ and $\hat{\pazocal{H}}$
To more clearly illustrate that in Fig. 20 (d) the transition in $\hat{\pazocal{H}}^{*}$ from chaotic to the fixed point dynamics near $b=-0.05$ happens via a bifurcation, in Fig. 26 we re-present Fig. 20 (d) from the perspective of the changes in the local minima of the $x_{3}$-variable. Also, to help further illustrate this bifurcation and the bifurcation from the limit cycle near $b=0.055$, in Figs. 27 and 28, we present snapshots of the RC’s dynamics at different $b$ values before and after these bifurcations take place.
Figure 27: Illustrating transition from $\hat{\pazocal{H}}^{*}$ to fixed point for decreasing $b$ as shown in Figs. 19 (d) and  26.
Figure 28: Illustrating transition from $\hat{\pazocal{L}}^{*}$ to limit cycle for decreasing $b$ as shown in Figs. 19 (d) and  26.
When looking at Figs. 26 and 27 together we see that there is no direct transition from chaotic attractor to fixed point, but rather the chaotic attractor that exists for $b=-0.056$ transitions to a limit cycle for $b=-0.055$. Prior to this transition, we see in Fig. 27 (c) that significant portions of the trajectory on this chaotic attractor are concentrated around where the limit cycle subsequently emerges. What is common among the limit cycle and chaotic attractor is that the trajectory around each attractor appears to contain a point nearby $(x_{1},x_{3})=(-6,26)$ and the state of the system spends a relatively large duration of time in the vicinity of this point. This suggests these attractors are of the saddle-focus homoclinic orbit type. Furthermore, it is this point (the saddle-focus) which becomes the stable fixed point for $b=-0.054$ shown in Fig. 27 (a).
In Fig. 28 we take a closer look at the transition in $\hat{\pazocal{L}}^{*}$ from Lorenz-like dynamics to the limit cycle near $b=0.06$ in Fig. 26. In Fig. 28 (a) we see that the state of the RC spends much larger durations of time in the left-hand wing of the attractor. The sharp corner in the trajectory nearby $(x_{1},x_{3})=(-1,12)$ indicates there is a saddle nearby, as flow converges towards this point along the $x_{3}$ direction and diverges away from this point along the $x_{1}$ direction. When decreasing $b$ from 0.0620.0620.062 in Fig. 28 (a) to 0.060.060.06 in Fig. 28 (b), it appears that the Lorenz-like attractor and saddle collide, which results in the right-hand wing disappearing and subsequently results in the period-4 limit cycle seen in Fig. 28 (b). We choose to show the change in the dynamics of this limit cycle from $b=0.06$ in Fig. 28 (b) to $b=0.04$ in Fig. 28 (b) in order to compliment the dynamics shown in Fig. 26, that the limit cycle continues to shrink until it becomes point-like and subsequently becomes the fixed point shown explicitly in Fig. 27 (a), which suggests that a Hopf bifurcation is responsible for this transition.
While we do not see regions of bistability between the GAs from the reconstructions of the different limit cycles in task (ii), the results presented in Figs. 20 (d) and 26 indicates that such a scenario is also possible. This leads us to make the following, more general, statement. If the RC fails to interpolate between reconstructed attractors via a sequence of continuous transitions then two options are possible: there is either a region of bistability between GAs that arise from the reconstructed attractors, or there is an UA that coexists with GAs or reconstructed attractors at different parameter settings.